OpenStudy (anonymous):

what is the figures showing an element of area and what is the area bounded by the curves y^2=4ax and x^2=4ay?

OpenStudy (tkhunny):

Can you find points of intersection between those two relations?

OpenStudy (anonymous):

0 and 0?

OpenStudy (tkhunny):

(0,0) is one point of intersection. You'll need another. Hint: There is symmetry about!

OpenStudy (anonymous):

Can you solve it for me? :)

OpenStudy (tkhunny):

Yes. When was the last time you factored a "Difference of Cubes"? You need to have your algebra up to speed!

OpenStudy (anonymous):

aww last 2 years?

OpenStudy (tkhunny):

Fair enough. It's time to do it again! Given these: y^2=4ax and x^2=4ay I recommend this solution: x^2=4ay Assuming everything is positive, there is no harm in squaring things. \(x^{4}=(4ay)^{2} = 16a^{2}y^{2}\) Are you bold enough to ponder x^4?! So far so good?

OpenStudy (anonymous):

y^2 = 4ax y = ±2√(ax) x^2 = 4ay y = x^2/4a A = ∫ 2√(ax) - x^2/(4a) dx [0,4a] = 4/3 * x √(ax) - x^3/(12a) [0,4a] = 4/3 * 4a * 2a - (64a^3)/(12a) = (32a^2)/3 - (16a^2)/3 = (16a^2)/3

OpenStudy (tkhunny):

Where did you get the limit of 4a?

OpenStudy (anonymous):

is it right? it is the same answer in the book

OpenStudy (tkhunny):

Answer my question. Magic won't help you on an exam.

OpenStudy (anonymous):

in the 2nd equation?

OpenStudy (tkhunny):

Nope. It's not there. You have it, but you have not proven it. How can you PROVE that 4a is the correct upper limit?